OpenStudy (anonymous):

A gas occupies a volume of 140 mL at 35.0 C and 97 kPa. What is the volume of the gas at STP

OpenStudy (chmvijay):

use gas equation

OpenStudy (chmvijay):

P1V1/T1=P2V2/T2

OpenStudy (anonymous):

ok how do u do that

OpenStudy (chmvijay):

V1=140 mL T1=35.0 C P1=97 kPa. V2=?,T2 =25 C ,P2=101.325 K Pa just mathematics find the V2

OpenStudy (chmvijay):

P1V1=nRT1 P2V2=nRT2 P1V1/T1 =nR P2V2/T2=nR hence we can write P1V1/T1=P2V2/T2

OpenStudy (chmvijay):

put them in above equation and solve for V2 did yo get me

OpenStudy (anonymous):

kinda